OpenStudy (anonymous):

b^2+8b+7-0 solve and check

OpenStudy (anonymous):

Solve: (b+7)(b+1)=0 b=-7, -1 Check: b=-1 =(-1)^2+8(-1)+7=0 =1-8+7=0 -7+7=0 0=0 b=-7 =(-7)^2+8(-7)+7=0 =49-56+7=0 -7+7=0 0=0 Done!